# Volumetric medical image segmentation through dual self-distillation in U-shaped networks

### v2

@@ 3.1 Experimental setup and implementation details
We attached our dual self-distillation (DSD) framework on several state-of-the-art U-shaped backbones and applied it on various benchmark datasets for 3D medical image segmentation tasks, specifically whole heart, brain tumor and Hippocampus segmentation.
@@
    nnU-Net [48]: The network consists of an input block to process the 3D input images, 3 downsample blocks, a bottleneck block at the bottom of the U-shaped structure and 4 upsample blocks to predict the ground-truth labels. For the input, downsample, bottleneck and upsample blocks, a convolution kernel of size (3,3,3) was used. For the input block, a stride of (1,1,1) was used whereas for the downsample and upsample blocks, a stride of (2,2,2) was used. For the MMWHS dataset, the number of CNN filters for the input block was 32 and the number of CNN filters for the downsample and upsample blocks were 64, 128 and 256. For the bottleneck layer, the number of filters were 320. For the MSD-BraTS and Hippocampus dataset, the number of CNN filters for the input block was 16 and the number of CNN filters for the downsample and upsample blocks were 32, 64 and 128. For the bottleneck layer, 256 filters were used.
@@
Figure 2: Qualitative comparison of an axial slice and 3D volumes, with ground-truth (GT) labels (on CTA) and predictions with (A) UNETR and (B) nnU-Net, highlighting the improved segmentations (shown by mean Dice score) with our proposed DSD framework.
@@ 3.2 Evaluation Metrics
    Dice score (Dice): The Dice similarity coefficient to measure the overlap between the ground-truth segmentation labels and the prediction from the network is given by:
@@
    where, $G_{p}$ and $Y_{p}$ denote the ground-truth labels and predicted output at voxel $p$. $N$ denotes the total number of voxels in the ground-truth and predicted output. A larger Dice score indicates a better overlap between ground-truth labels and predicted segmentations.
@@
    where, $G^{\prime}$ and $Y^{\prime}$ denote the ground-truth surface point set and predicted surface point set, respectively. A smaller HD indicates a better agreement between ground-truth and predicted surface point sets. The $95^{th}$ percentile of Hausdorff distance was used for all our experiments to reduce the effect of outliers.
3.3 Prediction with MMWHS dataset
Table 1 present the 5-fold cross validation results and summarizes the mean Dice score and HD95 of the 7 classes of cardiac substructures for the CT angiography (CTA) MMWHS dataset using UNETR and nnU-Net as the backbone along with the DSD framework. On average, across the 5-folds, when our proposed DSD framework is attached to the UNETR architecture, it outperforms the UNETR backbone with 2.7% increase in the mean Dice score and 3.22 mm decrease in mean HD95. On average, across the 5-folds, when our proposed DSD framework is attached to the nnU-Net architecture, it outperforms the nnU-Net backbone with 2.94% increase in the mean Dice score and 11.08 mm decrease in mean HD95. The improvements for both UNTER and nnU-Net with DSD attached to them, were achieved with negligible increase (less than 0.05%) in the number of training parameters and training time per epoch. In Table 1, we observe that the Dice score slightly drops for the MYO sub-structure when using nnU-Net with DSD. However, we also notice that there is a significant drop in the HD95 for MYO when using nnU-Net with DSD. Since, HD95 is a measure of boundary distance between the ground-truth and the predictions, it is clear that the basic nnU-Net (without DSD) has several isolated False predictions which leads to an increased boundary distance with the ground-truth labels. We also observe a significantly higher Dice score and lower HD95 for the complex and smaller sub-structures: Ascending Aorta (AA) and Pulmonary Artery (PA), which are clinically important and complex cardiac sub-structures that are harder to classify with a basic U-shaped network.
The results with the MMWHS dataset show that our DSD framework could still capture the intricate patterns in the CTA images and improve the segmentation performance over the U-shaped backbones, even when we have an anisotropic dataset and a high slice thickness along the axial direction. This confirms the renormalizabilty of our approach to any U-shaped network as well as to any type of datasets.
Additionally, as shown in Table 1, DSD also reduced the variance of the mean Dice score and mean HD95 across the 5-folds. A qualitative comparison (both on a 2D axial slice and 3D volume) between the ground-truth (GT) labels and prediction with UNETR backbone, UNETR with DSD, nnU-Net backbone and nnU-Net with DSD is shown in Fig. 2. We deliberately showed different poses for UNETR and nnU-Net, as the improvements after incorporating DSD to the backbones, could be better seen from those poses.
@@ 3.3 Prediction with MMWHS dataset
Figure 3: Qualitative comparison of an axial slice with ground-truth (GT) labels (on FLAIR MRI) and predictions from (A) UNETR, (B) nnU-Net and (C) Swin UNETR, highlighting the improvements in segmentation (shown by mean Dice score) with our proposed DSD framework.
@@ 3.4 Prediction with Brain Tumor dataset
We observe from Table 2 that the standard UNETR with DSD and nnU-Net with DSD achieves the best performance when compared with the state-of-the-art methods on the MSD-BraTS dataset [20, 55, 53, 54, 22, 48, 26]. When using the DSD framework with the U-shaped backbones on the MSD-BraTS dataset, we observed an increased mean and reduced variance of the Dice score of each class for all patients in the testing set as shown in Fig. 4. A qualitative comparison between the ground-truth (GT) labels and predictions with UNETR backbone, UNETR with DSD, nnU-Net backbone, nnU-Net with DSD, Swin UNETR backbone and Swin UNETR with DSD on a 2D slice (on FLAIR MRI) is shown in Fig. 3.
@@ 3.5 Prediction with Hippocampus dataset
The improvements for VNet, UNTER and nnU-Net with DSD attached to them, were achieved with negligible increase in the number of training parameters (less than 0.03%) and training time per epoch. When using the DSD framework with the U-shaped backbones on the MSD-Hippocampus dataset, we observed an increased mean and reduced variance of the Dice score and HD95 of each class for all patients in the testing set as shown in Table 3. A qualitative comparison between the ground-truth (GT) labels and predictions with VNet backbone, VNet with DSD, UNETR backbone, UNETR with DSD, nnU-Net backbone and nnU-Net with DSD on a 2D axial slice (on T1-w MRI) is shown in Fig. 5.
Figure 5: Qualitative comparison of an axial slice with ground-truth (GT) labels (on T1w MRI) and predictions from (A) VNET, (B) UNETR and (C) nnU-Net, highlighting the improvements in segmentation (shown by mean Dice score) with our proposed DSD framework.
@@ 3.7 Ablation Study
Table 5 presents the results of an ablation study with both UNETR and nnU-Net on one fold of the MMWHS validation set comparing the following components of our DSD framework: (i) Basic ($\eta=0$ and $\alpha_{1},\alpha_{2}=0$ in Eq. 5): This is the basic U-shaped backbone without any deep supervision and self-distillation. (ii) DS (Deep Supervision only: $\eta=1$ and $\alpha_{1},\alpha_{2}=0$ in Eq. 5): DSD in this case is reduced to deep supervision where the knowledge is distilled from the ground-truth labels to the deeper and shallower decoder layers of the U-shaped backbone. DS improves the Dice score by 2.17% over basic UNETR and 6.77% over basic nnU-Net. (iii) SDD (Self-Distillation between Decoders only: $\eta=1$, $\alpha_{1}=0$, $\alpha_{2}=1$ in Eq. 5): This shows the effect of self-distillation only between the decoder layers along with deep supervision in DSD. SDD improves the Dice score by 1.84% over UNETR with DS and 0.28% over nnU-Net with DS. (iv) SDE (Self-Distillation between Encoders only: $\eta=1$, $\alpha_{1}=1$, $\alpha_{2}=0$ in Eq. 5): This shows the effect of self-distillation only between the encoder layers along with deep supervision in DSD. SDE improves the Dice score by 1.61% over UNETR with DS and 0.68% over nnU-Net with DS. (v) DSD (Dual Self-Distillation: $\eta=1$, $\alpha_{1}=1$, $\alpha_{2}=1$ in Eq. 5): This shows the effect of our proposed dual self-distillation where knowledge distills from the deeper to the shallower layers both on the encoder and decoder side of the U-shaped backbone. DSD achieves the best performance with 2.64% and 2.36% increase in Dice score over UNETR with DS and nnU-Net with DS, respectively. We also performed quantitative studies on the choice of the temperature hyperparameter ($\tau$ in Eq. 2) for all our experiments. When we attached DSD to the nnU-Net architecture with $\tau$ = 1, 2, 3 and 4, we observed the Dice scores to be 85.37%, 85.89%, 86.10% and 86.07% respectively, on one-fold of the MMWHS validation set.
@@
We have performed additional ablation studies on one fold of the MMWHS validation set and the Hippocampus validation set as shown in Table 6. The Basic ((i) above) and DSD ((v) above) have the same settings as shown in Table 5. In the Mean Teacher experiment, the mean of all encoder outputs was used as teacher for all individual encoder outputs (students), and the mean of all decoder outputs was used as teacher for all individual decoder outputs (students) in the dual self-distillation process of the U-shaped networks. We observe that with this setting, there is also a significant drop in the dice score. For the Hippocampus dataset, we observed a 0.65% drop for UNETR and a 0.63% drop for the nnU-Net, compared to DSD. For the MMWHS dataset, we observed a 2.01% drop for UNETR and an 8.31% drop for the nnU-Net, compared to DSD. In this experiment, the representations for the teacher are smoothed-out due to the averaging operation (mean of all encoder or decoder representations) and hence this model could not surpass the performance of our DSD approach. In the Reversed Teacher experiment, we have reversed the direction of knowledge transfer on the encoder and decoder sides in our dual self-distillation framework, i.e., the knowledge was transferred in a top-down manner on the encoder side (from the shallowest encoder (Encoder 1) to the deeper encoders (Encoder 2, …, Encoder Z)) and in a bottom-up manner on the decoder side (from the shallowest decoder (Decoder Z) to the deeper decoders (Decoder 1, …, Decoder (Z-1))) [See Fig. 1]. We observe that with this setting, there is also a significant drop in the dice score which is due to the fact that we are reversing the direction of knowledge transfer between the encoders and decoders of the U-shaped network. For the Hippocampus dataset, we observed a 1.32% drop for UNETR and a 0.89% drop for the nnU-Net. For the MMWHS dataset, we observed a 3.7% drop for UNETR and an 8.95% drop for the nnU-Net. The rationale behind this experiment is to validate our design choice for DSD, i.e., to demonstrate that we are correctly distilling knowledge for both encoders and decoders in our DSD framework. The above experiments prove that the design choices made on both the encoder and decoder side in our DSD approach is appropriate.
@@
The original self-distillation implementation used the L2 loss between the feature maps of the teacher and student [30, 35]. In our ablation study, we also applied self-distillation from the feature maps of encoders and decoders. However, no noticeable improvement in performance was observed with this strategy, likely because we are performing pixel-level classification where the loss (Eq. 5) from each pixel is back-propagated to update the network weights, and hence the loss from feature maps become redundant. This is different from the image-level classification tasks [30, 35], where the loss from feature maps help improve the classification performance. Thus, feature map-based distillation was not included in the rest of our experiments.
@@ 3.8 Effect of parameters on network performance
Table 7 presents the effect of the number of trainable parameters on the performance of UNETR and nnU-Net on one fold of the MMWHS validation set. A small UNETR architecture with a feature size of 16 is compared with a large UNETR architecture with a feature size of 32. All the other settings were the same as the large UNETR, described in Section 3.1. When a small UNETR with less trainable parameters is used, applying the DSD to this small network improves the Dice score by 5.97% with only 0.014% increase in the number of trainable parameters and with negligible increase in training time per epoch. On the other hand, a large UNETR only leads to 3.21% increase in Dice score when compared with a small UNETR although there was a 9.7% increase in the number of trainable parameters with more training time per epoch. Finally, using DSD on the large UNETR improves the Dice score further by 4.81% over the large UNETR without DSD. Similarly, we compared a small nnU-Net with 16 and 256 filters for the input and bottleneck blocks and 32, 64 and 128 for both the downsample and upsample blocks. The other settings were the same as a large nnU-Net as described in Section 3.1. It should be noted that for the nnU-Net architecture, we only changed the initial setting for the number of filters for large and small networks such that the “Large Network” is initialized with much more filters while keeping all other hyperparameters the same. That is, the kernel sizes, strides and number of layers are all initialized the same for both “Small Network” and “Large Network”. Applying the DSD to a small nnU-Net improves the Dice score by 4.54% with only 0.14% increase in the number of parameters. In contrast, a large nnU-Net increases the Dice score by only 2.08% over the smaller nnU-Net although the number of trainable parameters increases by 193% with significant increase in training time per epoch. Finally, a large nnU-Net with DSD improves the Dice score further by 9.13% over the large nnU-Net without DSD, demonstrating that a nnU-Net initialized with a larger filter size cannot match the performance of a nnU-Net with DSD initialized with a smaller filter size.

### v3

@@ 3.1 Experimental setup and implementation details
We attached our DSD framework on several state-of-the-art U-shaped backbones and applied it on various benchmark datasets for 3D medical image segmentation tasks, specifically whole heart, brain tumor and Hippocampus segmentation.
@@
    nnU-Net [48]: The network consists of an input block to process the 3D input images, 3 downsample blocks, a bottleneck block at the bottom of the U-shaped structure and 4 upsample blocks to predict the GT Labels. For the input, downsample, bottleneck and upsample blocks, a convolution kernel of size (3,3,3) was used. For the input block, a stride of (1,1,1) was used whereas for the downsample and upsample blocks, a stride of (2,2,2) was used. For the MMWHS dataset, the number of CNN filters for the input block was 32 and the number of CNN filters for the downsample and upsample blocks were 64, 128 and 256. For the bottleneck layer, the number of filters were 320. For the MSD-BraTS and Hippocampus dataset, the number of CNN filters for the input block was 16 and the number of CNN filters for the downsample and upsample blocks were 32, 64 and 128. For the bottleneck layer, 256 filters were used.
@@
Figure 2: Qualitative comparison of an axial slice and 3D volumes, with GT Labels (on CTA) and predictions with (A) UNETR and (B) nnU-Net, highlighting the improved segmentations (shown by Dice score (%)) with our proposed DSD framework. The Dice score (%) is for the full 3D volume of a patient belonging to the validation set.
@@ 3.2 Evaluation Metrics
    Dice score (Dice): The Dice similarity coefficient to measure the overlap between the GT Labels and the prediction from the network is given by:
@@
    where, $G_{p}$ and $Y_{p}$ denote the GT Labels and predicted output at voxel $p$. $N$ denotes the total number of voxels in the GT Labels and the predicted output. A larger Dice score indicates a better overlap between GT Labels and predicted segmentations.
@@
    where, $G^{\prime}$ and $Y^{\prime}$ denote the GT surface point set and predicted surface point set, respectively. A smaller HD indicates a better agreement between GT and predicted surface point sets. The $95^{th}$ percentile of Hausdorff distance was used for all our experiments to reduce the effect of outliers.
@@
3.3 Prediction with MMWHS dataset
Table 1 present the 5-fold cross validation results and summarizes the mean Dice score and HD95 of the 7 classes of cardiac substructures for the CT angiography (CTA) MMWHS dataset using UNETR and nnU-Net as the backbone along with the DSD framework. On average, across the 5-folds, when our proposed DSD framework is attached to the UNETR architecture, it outperforms the UNETR backbone with 2.7% increase in the mean Dice score and 3.22 mm decrease in mean HD95. On average, across the 5-folds, when our proposed DSD framework is attached to the nnU-Net architecture, it outperforms the nnU-Net backbone with 2.94% increase in the mean Dice score and 11.08 mm decrease in mean HD95. The improvements for both UNTER and nnU-Net with DSD attached to them, were achieved with negligible increase (less than 0.05%) in the number of training parameters and training time per epoch. In Table 1, we observe that the Dice score slightly drops for the MYO sub-structure when using nnU-Net with DSD. However, we also notice that there is a significant improvement in the HD95 for MYO when using nnU-Net with DSD. Since, HD95 is a measure of boundary distance between the GT and the predictions, it is clear that the basic nnU-Net (without DSD) has several isolated False predictions which leads to an increased boundary distance with the GT Labels. We also observe a significantly higher Dice score and lower HD95 for the complex and smaller sub-structures: Ascending Aorta (AA) and Pulmonary Artery (PA), which are clinically important and complex cardiac sub-structures that are harder to classify with a basic U-shaped network.
The results with the MMWHS dataset show that our DSD framework could still capture the intricate patterns in the CTA images and improve the segmentation performance over the U-shaped backbones, even when we have an anisotropic dataset and a high slice thickness along the axial direction. This confirms that our approach could be applied to any U-shaped network and it works well with both isotropic and anisotropic datasets.
Additionally, as shown in Table 1, DSD also reduced the variance of the mean Dice score and mean HD95 across the 5-folds. A qualitative comparison (both on a 2D axial slice and 3D volume) between the GT Labels and prediction with UNETR backbone, UNETR with DSD, nnU-Net backbone and nnU-Net with DSD is shown in Fig. 2. We deliberately showed different poses for UNETR and nnU-Net, as the improvements after incorporating DSD to the backbones, could be better seen from those poses.
Figure 3: Qualitative comparison of an axial slice with GT Labels (on FLAIR MRI) and predictions from (A) UNETR, (B) nnU-Net and (C) Swin UNETR, highlighting the improvements in segmentation (shown by Dice score (%)) with our proposed DSD framework. The Dice score (%) is for the full 3D volume of a patient belonging to the testing set.
@@ 3.4 Prediction with Brain Tumor dataset
We observe from Table 2 that the standard UNETR with DSD and nnU-Net with DSD achieves the best performance when compared with the state-of-the-art methods on the MSD-BraTS dataset [20, 55, 53, 54, 22, 48, 26]. A qualitative comparison between the GT Labels and predictions with UNETR backbone, UNETR with DSD, nnU-Net backbone, nnU-Net with DSD, Swin UNETR backbone and Swin UNETR with DSD on a 2D slice (on FLAIR MRI) is shown in Fig. 3. When using the DSD framework with the U-shaped backbones on the MSD-BraTS dataset, we observed an increased mean and reduced variance of the Dice score of each class for all patients in the testing set as shown in Fig. 4.
@@ 3.5 Prediction with Hippocampus dataset
The improvements for VNet, UNTER and nnU-Net with DSD attached to them, were achieved with negligible increase in the number of training parameters (less than 0.03%) and training time per epoch. When using the DSD framework with the U-shaped backbones on the MSD-Hippocampus dataset, we observed an increased mean and reduced variance of the Dice score and HD95 of each class for all patients in the testing set as shown in Table 3. A qualitative comparison between the GT Labels and predictions with VNet backbone, VNet with DSD, UNETR backbone, UNETR with DSD, nnU-Net backbone and nnU-Net with DSD on a 2D axial slice (on T1-w MRI) is shown in Fig. 5. We would like to clarify that the Dice score reported in Fig. 5 pertains to the entire 3D volume, whereas only one 2D slice is depicted. In Row C, the improvement is distributed in smaller increments across many slices, leading to a sizable 9.08% overall gain. Conversely, for Rows A and B, we deliberately selected slices that visually accentuate the segmentation differences. Thus, even if Row C’s selected slice does not reveal a dramatic change, the collective effect across all slices in the volume contributes to the higher Dice score.
@@
Figure 5: Qualitative comparison of an axial slice with GT Labels (on T1w MRI) and predictions from (A) VNET, (B) UNETR and (C) nnU-Net, highlighting the improvements in segmentation (shown by Dice score (%)) with our proposed DSD framework. The Dice score (%) is for the full 3D volume of a patient belonging to the testing set.
@@ 3.7 Ablation Study
Table 5 presents the results of an ablation study with both UNETR and nnU-Net on one fold of the MMWHS validation set comparing the following components of our DSD framework: (i) Basic ($\eta=0$ and $\alpha_{1},\alpha_{2}=0$ in Eq. 5): This is the basic U-shaped backbone without any deep supervision and self-distillation. (ii) DS (Deep Supervision only: $\eta=1$ and $\alpha_{1},\alpha_{2}=0$ in Eq. 5): DSD in this case is reduced to deep supervision where the knowledge is distilled from the GT Labels to the deeper and shallower decoder layers of the U-shaped backbone. DS improves the Dice score by 2.17% over basic UNETR and 6.77% over basic nnU-Net. (iii) SDD (Self-Distillation between Decoders only: $\eta=1$, $\alpha_{1}=0$, $\alpha_{2}=1$ in Eq. 5): This shows the effect of self-distillation only between the decoder layers along with deep supervision in DSD. SDD improves the Dice score by 1.84% over UNETR with DS and 0.28% over nnU-Net with DS. (iv) SDE (Self-Distillation between Encoders only: $\eta=1$, $\alpha_{1}=1$, $\alpha_{2}=0$ in Eq. 5): This shows the effect of self-distillation only between the encoder layers along with deep supervision in DSD. SDE improves the Dice score by 1.61% over UNETR with DS and 0.68% over nnU-Net with DS. (v) DSD ($\eta=1$, $\alpha_{1}=1$, $\alpha_{2}=1$ in Eq. 5): This shows the effect of our proposed DSD where knowledge distills from the deeper to the shallower layers both on the encoder and decoder side of the U-shaped backbone. DSD achieves the best performance with 2.64% and 2.36% increase in Dice score over UNETR with DS and nnU-Net with DS, respectively. We also performed quantitative studies on the choice of the temperature hyperparameter ($\tau$ in Eq. 2) for all our experiments. When we attached DSD to the nnU-Net architecture with $\tau$ = 1, 2, 3 and 4, we observed the Dice scores to be 85.37%, 85.89%, 86.10% and 86.07% respectively, on one-fold of the MMWHS validation set.
The reason behind the improvement in performance with the different study settings in Table 5 is as follows: DS improves upon the Basic model by directly supervising intermediate decoder layers, forcing them to predict the segmentation mask and thereby mitigating vanishing gradients while promoting early learning of discriminative features. In SDD, the deepest decoder—with its refined semantic predictions—serves as a teacher to guide shallower decoders, ensuring consistency in feature representations and predictions. Similarly, in SDE, the deepest encoder, which captures rich contextual information, guides the shallower encoders to extract more robust features. DSD combines these strategies, simultaneously enhancing feature extraction in the encoder and prediction consistency in the decoder, leading to improved regularization, better gradient flow, and a more coherent network representation.
@@
We have also performed additional ablation studies on one fold of the MMWHS validation set and the Hippocampus validation set as shown in Table 6. The rationale behind these experiments is to validate our design choice for DSD, i.e., to demonstrate that we are correctly distilling knowledge for both encoders and decoders in our DSD framework. The Basic ((i) above) and DSD ((v) above) have the same settings as shown in Table 5. In the Mean Teacher experiment, the mean of all encoder outputs was used as teacher for all individual encoder outputs (students), and the mean of all decoder outputs was used as teacher for all individual decoder outputs (students) in the DSD process of the U-shaped networks. We observe that with this setting, there is a significant drop in the dice score. For the Hippocampus dataset, we observed a 0.65% drop for UNETR and a 0.63% drop for the nnU-Net, compared to DSD. For the MMWHS dataset, we observed a 2.01% drop for UNETR and an 8.31% drop for the nnU-Net, compared to DSD. In the Reversed Teacher experiment, we have reversed the direction of knowledge transfer on the encoder and decoder sides in our DSD framework, i.e., the knowledge was transferred in a top-down manner on the encoder side (from the shallowest encoder (Encoder 1) to the deeper encoders (Encoder 2, …, Encoder Z)) and in a bottom-up manner on the decoder side (from the shallowest decoder (Decoder Z) to the deeper decoders (Decoder 1, …, Decoder (Z-1))) [See Fig. 1]. For the Hippocampus dataset, we observed a 1.32% drop for UNETR and a 0.89% drop for the nnU-Net. For the MMWHS dataset, we observed a 3.7% drop for UNETR and an 8.95% drop for the nnU-Net.
The reason behind the poor performance of the Mean Teacher and Reversed Teacher method in Table 6 is as follows: In the Mean Teacher experiment, the representations for the teacher are smoothed-out due to the averaging operation (mean of all encoder or decoder representations), which produces a smoothed teacher signal that dilutes the refined, high-level details present in the deepest layers. This smoothing effect weakens the guidance provided to individual student layers, reducing the network’s ability to learn rich, context-specific features, and hence the Mean Teacher model underperformed our DSD approach. In the Reversed Teacher experiment, the performance drop is due to the fact that we are reversing the direction of knowledge transfer between the encoders and decoders of the U-shaped network. This reverses the natural information flow: in the encoder, shallower layers (with limited context) incorrectly guide deeper layers, and in the decoder, early-stage (less semantic) predictions teach later layers. This inversion undermines the hierarchical feature extraction and semantic refinement that are essential for accurate segmentation. The above experiments show that the design choices made on both the encoder and decoder side in our DSD approach are appropriate.
@@
The original self-distillation implementation used the L2 loss between the feature maps of the teacher and student [30, 35]. In our ablation study, we also applied the L2 loss to the self-distillation between the feature maps of encoders and decoders. However, no noticeable improvement in performance was observed with this strategy, likely because we are performing pixel-level classification where the loss (Eq. 5) from each pixel is back-propagated to update the network weights, and hence the loss from feature maps become redundant. This is different from the image-level classification tasks [30, 35], where the loss from feature maps help improve the classification performance. Thus, feature map-based distillation was not included in the rest of our experiments.
@@ 3.8 Effect of parameters on network performance
Table 7 presents the effect of the number of trainable parameters on the performance of UNETR and nnU-Net on one fold of the MMWHS validation set. A small UNETR architecture with a feature size of 16 is compared with a large UNETR architecture with a feature size of 32. All the other settings were the same as the large UNETR, described in Section 3.1. When a small UNETR with less trainable parameters is used, applying the DSD to this small network improves the Dice score by 5.97% with only 0.014% increase in the number of trainable parameters and with negligible increase in training time per epoch. On the other hand, a large UNETR only leads to 3.21% increase in Dice score when compared with a small UNETR although there was a 9.7% increase in the number of trainable parameters with more training time per epoch. Finally, using DSD on the large UNETR improves the Dice score further by 4.81% over the large UNETR without DSD. Similarly, we compared a small nnU-Net with 16 and 256 filters for the input and bottleneck blocks and 32, 64 and 128 for both the downsample and upsample blocks. The other settings were the same as a large nnU-Net as described in Section 3.1. It should be noted that for the nnU-Net architecture, we only changed the initial setting for the number of filters for large and small networks such that the “Large Network” is initialized with much more filters while keeping all other hyperparameters the same. That is, the kernel sizes, strides and number of layers are all initialized the same for both “Small Network” and “Large Network”. Applying the DSD to a small nnU-Net improves the Dice score by 4.54% with only 0.14% increase in the number of parameters. In contrast, a large nnU-Net increases the Dice score by only 2.08% over the smaller nnU-Net although the number of trainable parameters increases by 193% with significant increase in training time per epoch. Finally, a large nnU-Net with DSD improves the Dice score further by 9.13% over the large nnU-Net without DSD, demonstrating that a nnU-Net initialized with a larger number of filters cannot match the performance of a nnU-Net with DSD initialized with a smaller number of filters.